OpenStudy (anonymous):

When 4.0g of sulfur is burnt in excess oxygen, 40kJ of heat is evolved. How do you find the enthalpy change for the combustion of sulfur?

OpenStudy (anonymous):

S+O2 ---> SO2 (excess oxygen) 4.0 gm = 4/32 moles=1/8 moles if 1/8 moles give 40 kJ then 1 mole will give 40*8=320 kJ which is the standard enthalpy change for combustion of sulphur.